Nano Specialist is considering an upgrade project. The estimated cash flows from the upgrade project appear below. What is the project's payback period? Note that year 0 and year 1 cash flows are negative. (Answer in years, round to 2 places)

Year 0 cash flow = -79,000

Year 1 cash flow = -36,000

Year 2 cash flow = 24,000

Year 3 cash flow = 31,000

Year 4 cash flow = 27,000

Year 5 cash flow = 27,000

Year 6 cash flow = 26,000

Year 7 cash flow = 30,000
Answer #1

Cumulative cash flow for year 0 = -79,000

Cumulative cash flow for year 1 = -79,000 - 36,000 = -115,000

Cumulative cash flow for year 2 = -115,000 + 24,000 = -91,000

Cumulative cash flow for year 3 = -91,000 + 31,000 = -60,000

Cumulative cash flow for year 4 = -60,000 + 27,000 = -33,000

Cumulative cash flow for year 5 = -33,000 + 27,000 = -6,000

Cumulative cash flow for year 6 = -6,000 + 26,000 = 20,000

6,000 / 26,000 = 0.23

Payback period = 5 + 0.23 = 5.23 years